• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Keppel

    Log entry time 15:58:07 on November 2,2006

    Entry number 188755

    keyword=Shift Summary

    Started shift completing run 1908.

    Things went smoothly, until MCC called to take the beam for 10 minutes at ~8:20 AM. We ended run 1909.

    Beam restored, we started run 1910, and ran smoothly.

    ~10:10 AM Hall C took out the 1/2 wave plate.

    We kept running until ~10:30 AM, when MCC took the beam for a spot move. Ended with run 1912.

    Since there was no beam in the hall, Jack Segal made an opportunistic access to work on the NMR and Alexandre Camsonne made one as well to work on the Compton.

    Hall C had the half wave plate put back in after their Moller work.

    The spot move transitioned into Hall C Moller running (which our beam interferes with) and then beam studies, and then problems, so the beam did not come back until ~3:30 PM. We started run 1913.

    This was the last production run for our shift...Sirish Nanda came in to begin Compton studies at low current at ~4 PM.